# Construction of the log-convex minorant of a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}_{0}^{d}}$

### 20 Jan 2024 v1

@@ 1. Introduction
Mandelbrojt proved in [M, Chap. I] (see also [K]) that if $\lim_{p\to+\infty}M_{p}^{1/p}=+\infty$ then
@@
We refer also to the recent work [S] where this construction has been studied again in detail, some technical ambiguities have been solved and non-standard cases have been studied as well; see Section 3.
However, to the best of our knowledge this condition has never been generalized to the $d$-dimensional anisotropic case ($d>1$), and the reason is that the classical coordinate-wise logarithmic convexity condition (5.8) for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is not sufficient to obtain the analogous of (1.1) for $M_{\alpha}$, as explained in Remark 5.3 (here $e_{j}$ is the $j$-th element of the canonical basis of $\mathbb{R}^{d}$ with all entries equal to 0 except the $j$-th entry equal to 1). The reason is that this is a convexity condition on each variable separately and not on the globality of its variables. Assuming the stronger condition that $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is log-convex on the globality of its variables $\alpha\in\mathbb{N}_{0}^{d}$ (see Definition 5.1), we extend (1.1) to $\alpha\in\mathbb{N}_{0}^{d}$ instead of $p\in\mathbb{N}_{0}$ (see Theorem 5.2).
To obtain this result we construct in Sections 2-4 the (optimal) convex minorant of a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ (then $a_{\alpha}=\log M_{\alpha}$ in Section 5). The idea, in the one variable case, takes inspiration from the convex regularization of sequences of Mandelbrojt in [M], which was quite complicated and difficult to export to the more dimensional case. Our construction is made by taking the supremum of hyperplanes approaching from below the given sequence and leads to the notion of convexity for a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ in the sense that $a_{\alpha}=F(\alpha)$ for a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$. This condition gives the suitable notion of logarithmic convexity for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ in order to write it in terms of its associated function as in (5.7).
As an application of the above results, in Section 6 we characterize the condition on the matrix weights $\mathcal{M}$ and $\mathcal{N}$ in order to have a continuous inclusion between the spaces $\mathcal{S}_{\{\mathcal{M}\}}/\mathcal{S}_{(\mathcal{M})}$, $\mathcal{S}_{\{\mathcal{N}\}}/\mathcal{S}_{(\mathcal{N})}$ of rapidly decreasing ultradifferentiable functions in the matrix weighted setting, both in the Roumieu and Beurling cases.
@@ 3. Geometric construction of the convex minorant sequence in the one-dimensional case
The geometric construction in the one-dimensional case can also be found in [S]; we refer to this work and in particular to [S, Sect. 3.3] concerning the comparison with the classical results from [M, Chap. I]. Here we revisit the construction in the spirit of Section 2 since it is required in the induction argument for the higher-dimensional situation in Section 4.
@@ 4. Convex minorant sequence in the ${d}$-dimensional case
for some $k^{*}\in\mathbb{R}^{d}$ and $c^{*}=y^{0}-\langle k^{*},x^{0}\rangle$, that leaves the whole set $G_{g}^{+}$ on the same side of the hyperplane. Note that $x^{0}\in(0,+\infty)^{d}$ avoids “vertical” hyperplanes of the form $x_{j}=c$, because of the convexity of $g$.
@@
In order to prove the opposite inequality we remark that if $f\in\mathcal{L}_{S}$, i.e.
$f(x)=\langle k,x\rangle+c,\quad k\in\mathbb{R}^{d},\,c\in\mathbb{R},$
(4.5) with $\langle k,\alpha\rangle+c\leq a_{\alpha},\quad\forall\alpha\in\mathbb{N}_{0}^{d},$
then
$f(x^{\prime},0)=\langle k^{\prime},x^{\prime}\rangle+c\in\mathcal{L}_{S}^{0}$
since $\langle k^{\prime},\alpha^{\prime}\rangle+c\leq a_{(\alpha^{\prime},0)}$ by (4.5).
Therefore
$F^{0}(x^{\prime})=\sup_{h\in\mathcal{L}_{S}^{0}}h(x^{\prime})\geq\sup_{f\in\mathcal{L}_{S}}f(x^{\prime},0)=F(x^{\prime},0)$
and the equality (4.4) is proved. This means that we have reduced the problem to prove that $F^{0}(x^{\prime})$ is the maximum convex function $g:[0,+\infty)^{d-1}\to\mathbb{R}$ such that
@@ 5. Construction of the log-convex minorant
Condition (5.1) ensures that $\omega_{\mathbf{M}}(t)<+\infty$ for all $t\in\mathbb{R}^{d}$ (see [M, Chap. I] or [BJOS1, Rem. 1]).
@@
Note that if $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is a log-convex sequence, i.e. $\log M_{\alpha}=F(\alpha)$ for a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$, then
@@
This condition is clearly weaker than the condition of logarithmic convexity given in Definition 5.1 since there are functions which are coordinate-wise convex but not convex as functions of more variables. In particular, from (5.5) we have that (5.8) is not sufficient to obtain (5.7); for an explicit example, see Example 5.4 below. Clearly in the one-dimensional case the two notions of log-convexity coincide, and (5.7) was already known (see [K, Prop. 3.2]).
@@ 6. Characterization of inclusion relations of spaces of rapidly decreasing ultradifferentiable functions
Let us first recall the notion of weight matrices (anisotropic framework) as considered in [BJOS1, Sec. 3].
@@
endowed with the inductive limit topology in the Roumieu case and the projective limit topology in the Beurling case (see [BJOS1, Sec. 3]).
@@
We shall also need the following conditions for a weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ (see [BJOS1]; cf. also [L]), in the Roumieu setting
@@
Note that (6.1) for $\beta=0$ implies that $\mathbf{M}^{(\kappa)}$ satisfies (5.1) for some $\kappa>0$, and hence for all $\kappa^{\prime}\geq\kappa$. Similarly (6.4) implies that $\mathbf{M}^{(\lambda)}$ satisfies (5.1) for all $\lambda>0$ (see [BJOS1, Rem. 3]). Since condition (5.1) ensures that the associate weight function $\omega_{\mathbf{M}}$ is finite, the above remarks are essential to recall, from [BJOS1, Thm. 1], that if the weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ satisfies (6.1) and (6.2) then $\mathcal{S}_{\{\mathcal{M}\}}$ is isomorphic to the sequence space
@@
It will also be useful, in the following, to write the sequence spaces (6.7)-(6.8) as follows (see [BJOS1]):
@@
$\Lambda_{(\mathcal{M})}=\{\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}:\ \forall j\in\mathbb{N}\,\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(1/j)}}(j\alpha^{1/2})}<+\infty\}.$
@@
The above remarks explain why conditions (6.3) and (6.6), that we shall need in Theorem 6.1, are not required in the one-dimensional/isotropic case (see [BJOS2]).
@@
Acknowledgments. Boiti and Oliaro were partially supported by the INdAM - GNAMPA Project 2023 “Analisi di Fourier e Analisi Tempo-Frequenza di Spazi Funzionali ed Operatori”. Boiti was partially supported by the Projects FAR 2020, FAR 2021, FAR 2022, FIRD 2022 and FAR 2023 (University of Ferrara). Jornet is partially supported by the project PID2020-119457GB-100 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. Schindl is supported by FWF (Austrian Science fund) project 10.55776/P33417.

### 12 Apr 2024 v2

@@ 1. Introduction
Mandelbrojt proved in [12, Chap. I] (see also [10]) that if $\lim_{p\to+\infty}M_{p}^{1/p}=+\infty$ then
@@
We refer also to the recent work [16] where this construction has been studied again in detail, some technical ambiguities have been solved and non-standard cases have been studied as well; see Section 3.
However, to the best of our knowledge this condition has never been generalized to the $d$-dimensional anisotropic case ($d>1$), and the reason is that the classical coordinate-wise logarithmic convexity condition (5.8) for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is not sufficient to obtain the analogous of (1.1) for $M_{\alpha}$, as explained in Remark 5.3. The reason is that this is a convexity condition on each variable separately and not on the globality of its variables. Assuming the stronger condition that $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is log-convex on the globality of its variables $\alpha\in\mathbb{N}_{0}^{d}$ (see Definition 5.1), we extend (1.1) to $\alpha\in\mathbb{N}_{0}^{d}$ instead of $p\in\mathbb{N}_{0}$ (see Theorem 5.2).
To obtain this result we construct in Sections 2-4 the (optimal) convex minorant of a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ (then $a_{\alpha}=\log M_{\alpha}$ in Section 5). The idea, in the one variable case, takes inspiration from the convex regularization of sequences of Mandelbrojt in [12], which was quite complicated and difficult to export to the more dimensional case. Our construction is made by taking the supremum of hyperplanes approaching from below the given sequence and leads to the notion of convexity for a sequence $\{a_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ in the sense that $a_{\alpha}=F(\alpha)$ for a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$. This condition gives the suitable notion of logarithmic convexity for a sequence $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ in order to write it in terms of its associated function as in (5.7).
This result is a very useful tool for working in the anisotropic setting. In Section 6, indeed, we obtain some applications about inclusion of spaces of rapidly decreasing ultradifferentiable functions in the matrix weighted anisotropic setting, where Theorem 5.2 is crucial (see Remark 6.2 and compare with the isotropic case in [3]). In particular, we characterize the conditions on the matrix weights $\mathcal{M}$ and $\mathcal{N}$ in order to have a continuous inclusion between the spaces $\mathcal{S}_{\{\mathcal{M}\}}/\mathcal{S}_{(\mathcal{M})}$, $\mathcal{S}_{\{\mathcal{N}\}}/\mathcal{S}_{(\mathcal{N})}$, both in the Roumieu and Beurling cases. The advantages of the matrix weighted setting was already enlightened in [13], in order to treat at the same time classes in the sense of Komatsu [10] (estimates of the derivatives with a sequence) and in the sense of Braun, Meise and Taylor [5] (estimates of the derivatives via a weight function). Since then several papers using weight matrices have been published. We mention, for instance, [2, 7, 8, 9], and the references therein.
We refer to [4] to compare classes of ultradifferentiable functions as defined by Komatsu [10] and as defined by Braun, Meise and Taylor [5], and we point out the recent papers [17, 14, 15, 1, 6] for other different interesting results in anisotropic Gelfand-Shilov spaces.
@@ 3. Geometric construction of the convex minorant sequence in the one-dimensional case
The geometric construction in the one-dimensional case can also be found in [16]; we refer to this work and in particular to [16, Sect. 3.3] concerning the comparison with the classical results from [12, Chap. I]. Here we revisit the construction in the spirit of Section 2 since it is required in the induction argument for the higher-dimensional situation in Section 4.
@@ 4. Convex minorant sequence in the ${d}$-dimensional case
for some $k^{*}\in\mathbb{R}^{d}$ and $c^{*}=y^{0}-\langle k^{*},x^{0}\rangle$, that leaves the whole set $G_{g}^{+}$ on the same side of the hyperplane. Note that $x^{0}\in(0,+\infty)^{d}$ avoids “vertical” hyperplanes.
@@
In order to prove the opposite inequality, let us now fix $f\in\mathcal{L}_{S}^{0}$. Then
$f(x^{\prime})=\langle k^{\prime},x^{\prime}\rangle+c,\quad k^{\prime}\in\mathbb{R}^{d-1},\,c\in\mathbb{R},$
and
(4.5) $f(\alpha^{\prime})\leq a_{(\alpha^{\prime},0)}\quad\forall\alpha^{\prime}\in\mathbb{N}_{0}^{d-1}.$
We claim that there exists $\Lambda\in\mathbb{R}$ such that
(4.6) $\tilde{f}(x):=f(x^{\prime})+\Lambda x_{d}\in\mathcal{L}_{s}$
for $x=(x^{\prime},x_{d})\in[0,+\infty)^{d}$.
Indeed, as in Lemma 2.3, from assumption (ii) and $|\langle k^{\prime},\alpha^{\prime}\rangle+c|\leq\|k^{\prime}\|\cdot|\alpha|+|c|$ we get that
$\langle k^{\prime},\alpha^{\prime}\rangle+c>a_{\alpha},$
for at most a finite number of points $\alpha=(\alpha^{\prime},\alpha_{d})\in\mathbb{N}_{0}^{d}$, with $\alpha_{d}\geq 1$ (because of (4.5) for $\alpha_{d}=0$).
There exists then
$L:=\min_{\alpha\in\mathbb{N}_{0}^{d}}[a_{\alpha}-(\langle k^{\prime},\alpha^{\prime}\rangle+c)].$
Defining
$\Lambda:=\min(0,L),$
we have that $\Lambda\leq 0$ and hence for all $\alpha=(\alpha^{\prime},\alpha_{d})\in\mathbb{N}_{0}^{d}$ we have two cases:
if $\alpha_{d}=0$, then by (4.5)
$\langle k^{\prime},\alpha^{\prime}\rangle+c+\Lambda\alpha_{d}\leq a_{(\alpha^{\prime},0)};$
if $\alpha_{d}\geq 1$, then by definition of $L$
$\langle k^{\prime},\alpha^{\prime}\rangle+c+\Lambda\alpha_{d}\leq\langle k^{\prime},\alpha^{\prime}\rangle+c+\Lambda\leq\langle k^{\prime},\alpha^{\prime}\rangle+c+L\leq a_{\alpha}.$
The two cases above prove (4.6).
Setting now
$\tilde{k}:=(k^{\prime},\Lambda)\in\mathbb{R}^{d}\quad\mbox{and}\quad\tilde{f}(x):=\langle\tilde{k},x\rangle+c,$
we have that $\tilde{f}\in\mathcal{L}_{S}$ and $\tilde{f}(x^{\prime},0)=f(x^{\prime})$.
It follows that
$F^{0}(x^{\prime})$ $=\sup_{f\in\mathcal{L}_{S}^{0}}f(x^{\prime})=\sup_{f\in\mathcal{L}_{S}^{0}}\tilde{f}(x^{\prime},0)$
$\leq\sup_{f\in\mathcal{L}_{S}}f(x^{\prime},0)=F(x^{\prime},0),$
and the equality (4.4) is therefore proved.
This means that we have reduced the problem to prove that $F^{0}(x^{\prime})$ is the maximum convex function $g:[0,+\infty)^{d-1}\to\mathbb{R}$ such that
@@ 5. Construction of the log-convex minorant
Condition (5.1) ensures that $\omega_{\mathbf{M}}(t)<+\infty$ for all $t\in\mathbb{R}^{d}$ (see [12, Chap. I] or [2, Rem. 1]).
@@
Note that if $\{M_{\alpha}\}_{\alpha\in\mathbb{N}^{d}_{0}}$ is a log-convex sequence, i.e. $\log M_{\alpha}=F(\alpha)$ for a convex function $F:[0,+\infty)^{d}\to\mathbb{R}$, and $e_{j}$ denotes the $j$-th element of the canonical basis of $\mathbb{R}^{d}$ with all entries equal to 0 except the $j$-th entry equal to 1, then
@@
This condition is clearly weaker than the condition of logarithmic convexity given in Definition 5.1 since there are functions which are coordinate-wise convex but not convex as functions of more variables. In particular, from (5.5) we have that (5.8) is not sufficient to obtain (5.7); for an explicit example, see Example 5.4 below. Clearly in the one-dimensional case the two notions of log-convexity coincide, and (5.7) was already known (see [10, Prop. 3.2]).
@@ 6. Characterization of inclusion relations of spaces of rapidly decreasing ultradifferentiable functions
Let us first recall the notion of weight matrices (anisotropic framework) as considered in [2, Sec. 3].
@@
endowed with the inductive limit topology in the Roumieu case and the projective limit topology in the Beurling case (see [2, Sec. 3]).
@@
We shall also need the following conditions for a weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ (see [2]; cf. also [11]), in the Roumieu setting
@@
Note that (6.1) for $\beta=0$ implies that $\mathbf{M}^{(\kappa)}$ satisfies (5.1) for some $\kappa>0$, and hence for all $\kappa^{\prime}\geq\kappa$. Similarly (6.4) implies that $\mathbf{M}^{(\lambda)}$ satisfies (5.1) for all $\lambda>0$ (see [2, Rem. 3]). Since condition (5.1) ensures that the associate weight function $\omega_{\mathbf{M}}$ is finite, the above remarks are essential to recall, from [2, Thm. 1], that if the weight matrix =$\mathcal{M}=\{(\mathbf{M}^{(\lambda)})_{\lambda>0}\}$ satisfies (6.1) and (6.2) then $\mathcal{S}_{\{\mathcal{M}\}}$ is isomorphic to the sequence space
@@
It will also be useful, in the following, to write the sequence spaces (6.7)-(6.8) as follows (see [2]):
@@
(6.10) $\Lambda_{(\mathcal{M})}=\{\mathbf{c}=(c_{\alpha})_{\alpha\in\mathbb{N}_{0}^{d}}\in\mathbb{C}^{\mathbb{N}_{0}^{d}}:\ \forall j\in\mathbb{N}\,\sup_{\alpha\in\mathbb{N}_{0}^{d}}|c_{\alpha}|e^{\omega_{\mathbf{M}^{(1/j)}}(j\alpha^{1/2})}<+\infty\}.$
@@
The above remarks explain why conditions (6.3) and (6.6), that we shall need in Theorem 6.1, are not required in the one-dimensional/isotropic case (see [3]).
@@
Remark 6.2.
In Theorem 6.1 we used Theorem 5.2 to characterise the inclusion relations of the spaces for any dimension $d$, which was not possible in the analogous results [3, Theorems 4.4, 4.5, 4.6], where we needed $d=1$ in one implication since formula (5.7) was not available for the general anisotropic case.
Acknowledgments. The authors are sincerely grateful to Prof. Andreas Debrouwere for pointing out that the argument in the first version of the proof of Theorem 4.1 was incomplete. Boiti and Oliaro were partially supported by the INdAM - GNAMPA Project 2023 “Analisi di Fourier e Analisi Tempo-Frequenza di Spazi Funzionali ed Operatori”, CUP_ E53C22001930001. Boiti was partially supported by the Projects FAR 2020, FAR 2021, FAR 2022, FIRD 2022 and FAR 2023 (University of Ferrara). Jornet is partially supported by the project PID2020-119457GB-100 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. Schindl is supported by FWF (Austrian Science fund) project 10.55776/P33417.
Statements and Declarations:
  • Conflict of interest. The authors declare that there is no conflict of interest.
  • Data availability statements. The manuscript has no associated data.